OpenStudy (anonymous):

Find the Indicate Outputs: f(x)=4x-3x f(0)= f(-1)= f(2)=

OpenStudy (saifoo.khan):

is its 4x^2 ??

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

1) 0 2) -1 3) 2

OpenStudy (anonymous):

got it. thanks

OpenStudy (saifoo.khan):

medal?

OpenStudy (anonymous):

f(x)=4x-3x f(0)=0 f(-1)=-4+3=-1 f(2)=8-6=2

OpenStudy (saifoo.khan):

bus kar naro, phy ka ques ka ans day!